# Group inverses of matrices of simple symmetric directed graphs

### v1

@@ 2. Graphical Description of $A^{\#}$ for a Matrix with Simple Symmetric Digraph
Let $A=(a_{ij})$ be an $n\times n$ real matrix such that $D(A)\in\mathbb{D}$. Let $\{i_{1},i_{2},\ldots,i_{s+1}\}$ be a set of all pendant vertices with a common neighbour $q$. Then,
@@
In order to prove the main result, we need three lemmata, which we prove now.
@@
Let $A=(a_{ij})$ be an $n\times n$ real matrix such that $D(A)\in\mathbb{D}$. Let $D(A)$ be a strongly connected graph and $\Delta_{A}\neq 0$. Let $B=(b_{ij})$ be the matrix given by $b_{ij}=\frac{\mu_{ij}}{\Delta_{A}},1\leq i,j\leq n$. Then, $AB=BA$.
@@
Now, we consider the fact $i\neq j$ and prove 2.4 by considering four cases.
@@
Subcase (iii.1): $i$ is adjacent to $j$. then, the left hand side of 2.4 is zero and the right hand side equal to
@@
Subcase (iii.2.a): $q$ is adjacent to $j$. Since non-pendant cycles do not belong to any maximum matching, again, the left hand side of 2.4 is zero, and the right hand side can be split as 2.5. Let $\mathbb{M}_{j_{l}}(i,q)\subseteq\mathbb{M}(i,q)$ be the set of maximum matching containing pendant cycle $(j,j_{l},j)$. Then $\mathbb{M}(i,q)=\cup_{l=1}^{s}\mathbb{M}_{j_{l}}(i,q)$, a mutually disjoint union. Now, the second sum of 2.5 is
@@
So, the right hand side of 2.4 is also zero.
Subcase (iii.2.b): $q$ is not adjacent to $j$. Since $q$ and $j$ both are non-pendant vertices, the left hand side of equation 2.4 is zero. After splitting the right hand side as 2.5, in the first sum, the length of any cycle chain from vertex $i$ to $j_{l}$ must be at least four for any $l\in\{1,2,\ldots,s\}$ and so, $\mu_{ij_{j}}$ is zero. Next, as $m$ is a non-pendant vertex and since it is not adjacent to $i$, the second sum is also zero, showing that the right hand side equals zero.
@@
For a pendant vertex $i$, let $\{i_{1},i_{2},\ldots,i_{s}\}$ be the set of all pendant vertices other than $i$, which have a common neighbour.
Case (i): $(i,j)$ is an edge.
Subcase (i.1): $i$ is a non-pendant vertex. By Corollary 2.11, $(AB)_{ii}=1$ and $(AB)_{ik}=0$ for any $k\neq i$. So, $c=a_{ij}$.
Subcase (i.2): $i$ is a pendant vertex. Then, $j$ is a non-pendant vertex adjacent to $i$. By Remark 2.9, $\mathbb{M}=\mathbb{M}(i)\cup_{m=1}^{s}\mathbb{M}(i_{m})$, a mutually disjoint union. By Corollary 2.11, $c_{i}=\frac{\sum_{M\in\mathbb{M}(i)}\eta(M)}{\Delta_{A}}a_{ij}$ and
$\tilde{c_{i}}$ $=\sum_{m=1}^{s}(AB)_{ii_{m}}a_{i_{m}j}$
$=\sum_{m=1}^{s}\Big{(}\frac{a_{ji_{m}}\mu_{ij}}{\Delta_{A}}\Big{)}a_{i_{m}j}$
$=\frac{1}{\Delta_{A}}\sum_{m=1}^{s}a_{ji_{m}}a_{i_{m}j}\Big{(}\beta_{ij}\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\Big{)}$
$=\frac{1}{\Delta_{A}}\sum_{m=1}^{s}a_{ji_{m}}a_{i_{m}j}\Big{(}a_{ij}\sum_{M\in\mathbb{M}(i_{m},j)}\beta_{\overline{i_{m},j}}(M)\Big{)}$
$=\frac{a_{ij}}{\Delta_{A}}\sum_{m=1}^{s}\Big{(}a_{ji_{m}}a_{i_{m}j}\sum_{M\in\mathbb{M}(i_{m},j)}\beta_{\overline{i_{m},j}}(M)\Big{)}$
$=\frac{a_{ij}}{\Delta_{A}}\sum_{m=1}^{s}\sum_{M\in\mathbb{M}(i_{m})}\eta(M).$
Thus, $c=\frac{a_{ij}}{\Delta_{A}}\Big{(}\sum_{M\in\mathbb{M}(i)}\eta(M)+\sum_{m=1}^{s}\sum_{M\in\mathbb{M}(i_{m})}\eta(M)\Big{)}=\frac{a_{ij}}{\Delta_{A}}\cdot\Delta_{A}=a_{ij}$.
Case (ii): $(i,j)$ is not an edge. Then, $a_{ij}=0$. So, $c=\tilde{c_{i}}$.
Subcase (ii.1): $i$ is a non-pendant vertex. Then, by Remark 2.12, $(AB)_{ik}$ is zero for any $k\neq i$. So, $c=\tilde{c_{i}}=0$.
Subcase (ii.2): $i$ is a pendant vertex. Then, again by Remark 2.12, $\tilde{c_{i}}=\sum_{m=1}^{s}(AB)_{ii_{m}}a_{i_{m}j}$. Since $i$ is not adjacent to $j$, $i_{m}$ is also not adjacent to $j$ for any $m\in\{1,2,\ldots,s\}$. Thus, $c=\tilde{c_{i}}=0$, completing the proof. ∎
@@
Case (i): $i,j$ are maximally matchable. Then, by Proposition 2.5, the length of alternating cycle chain from $i$ to $j$ is at most three.
Subcase (i.1): The length of alternating cycle chain from $i$ to $j$ is three. Then, $i$ and $j$ are both must be pendant vertex. Let $\{i_{1},i_{2},\ldots,i_{s}\}$ be the set of all pendant vertices other than $i$, which have a common neighbour $q$. Let the unique alternating cycle chain from $i$ to $j$ be $((i,q,i),(q,z,q),(z,j,z))$. Again, by Corollary 2.11, $b_{i}=\frac{\sum_{M\in\mathbb{M}(i)}\eta(M)}{\Delta_{A}}\mu_{ij}$ and
$\tilde{b_{i}}$ $=\sum_{m=1}^{s}(AB)_{ii_{m}}\mu_{i_{m}j}$
$=\sum_{m=1}^{s}\Big{(}\frac{a_{qi_{m}}\mu_{iq}}{\Delta_{A}}\Big{)}\mu_{i_{m}j}$
$=\frac{1}{\Delta_{A}}\sum_{m=1}^{s}a_{qi_{m}}\Big{(}\beta_{iq}\sum_{M\in\mathbb{M}(i,q)}\beta_{\overline{i,q}}(M)\Big{)}\Big{(}\beta_{i_{m}j}\sum_{M\in\mathbb{M}(i_{m},j)}\beta_{\overline{i_{m},j}}(M)\Big{)}$
$=\frac{1}{\Delta_{A}}\sum_{m=1}^{s}a_{qi_{m}}\Big{(}a_{iq}\sum_{M\in\mathbb{M}(i_{m},q)}\beta_{\overline{i_{m},q}}(M)\Big{)}\Big{(}-a_{i_{m}q}a_{qz}a_{zj}\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\Big{)}$
$=\frac{1}{\Delta_{A}}\sum_{m=1}^{s}\Big{(}a_{i_{m}q}a_{qi_{m}}\sum_{M\in\mathbb{M}(i_{m},q)}\beta_{\overline{i_{m},q}}(M)\Big{)}\Big{(}-a_{iq}a_{qz}a_{zj}\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\Big{)}$
$=\frac{1}{\Delta_{A}}\sum_{m=1}^{s}\Big{(}\sum_{M\in\mathbb{M}(i_{m})}\eta(M)\Big{)}\Big{(}\beta_{ij}\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\Big{)}$
$=\frac{\mu_{ij}}{\Delta_{A}}\sum_{m=1}^{s}\sum_{M\in\mathbb{M}(i_{m})}\eta(M).$
Thus, by Remark 2.9, $b=\frac{\mu_{ij}}{\Delta_{A}}\Big{(}\sum_{M\in\mathbb{M}(i)}\eta(M)+\sum_{m=1}^{s}\sum_{M\in\mathbb{M}(i_{m})}\eta(M)\Big{)}=\frac{\mu_{ij}}{\Delta_{A}}\cdot\Delta_{A}=\mu_{ij}$.
Subcase (i.2): The length of alternating cycle chain from $i$ to $j$ is one. Then, $i$ is adjacent to $j$.
Subcase (i.2.a): $i$ is a non-pendant vertex. Then, $(AB)_{ii}=1$ and $(AB)_{ik}=0$ for any $k\neq i$. So, $b=\mu_{ij}$.
Subcase (i.2.b): $i$ is a pendant vertex. Let $\{i_{1},i_{2},\ldots,i_{s}\}$ be the set of all pendant vertices other than $i$, which have a common neighbour $j$. Then, by Remark 2.9 and 2.12, $b_{i}=\frac{\sum_{M\in\mathbb{M}(i)}\eta(M)}{\Delta_{A}}\mu_{ij}$ and
$\tilde{b_{i}}$ $=\sum_{m=1}^{s}(AB)_{ii_{m}}\mu_{i_{m}j}$
$=\sum_{m=1}^{s}\Big{(}\frac{a_{ji_{m}}\mu_{ij}}{\Delta_{A}}\Big{)}\mu_{i_{m}j}$
$=\frac{\mu_{ij}}{\Delta_{A}}\sum_{m=1}^{s}a_{ji_{m}}\Big{(}\beta_{i_{m}j}\sum_{M\in\mathbb{M}(i_{m},j)}\beta_{\overline{i_{m},j}}(M)\Big{)}$
$=\frac{\mu_{ij}}{\Delta_{A}}\sum_{m=1}^{s}\Big{(}a_{ji_{m}}a_{i_{m}j}\sum_{M\in\mathbb{M}(i_{m},j)}\beta_{\overline{i_{m},j}}(M)\Big{)}$
$=\frac{\mu_{ij}}{\Delta_{A}}\sum_{m=1}^{s}\sum_{M\in\mathbb{M}(i_{m})}\eta(M).$
Thus, by Remark 2.9, $b=\frac{\mu_{ij}}{\Delta_{A}}\Big{(}\sum_{M\in\mathbb{M}(i)}\eta(M)+\sum_{m=1}^{s}\sum_{M\in\mathbb{M}(i_{m})}\eta(M)\Big{)}=\frac{\mu_{ij}}{\Delta_{A}}\cdot\Delta_{A}=\mu_{ij}$.
Case (ii): $i,j$ are not maximally matchable. Then, $\mu_{ij}=0$ and $b=\tilde{b_{i}}$.
Subcase (ii.1): $i$ is pendant vertex. Let $\{i_{1},i_{2},\ldots,i_{s}\}$ be the set of all pendant vertices other than $i$, which have a common neighbour. Then, by Remark 2.12, $\tilde{b_{i}}=\sum_{m=1}^{s}(AB)_{ii_{m}}\mu_{i_{m}j}$. Since $\{i,i_{1},i_{2},\ldots,i_{s}\}$ be a set of all pendant vertices with a common neighbour and $i,j$ are not maximally matchable, $i_{m},j$ are also not maximally matchable for any $m\in\{1,2,\ldots,s\}$. So, $\mu_{i_{m}j}=0$ for all $m\in\{1,2,\ldots,s\}$. Thus, $b=\tilde{b_{i}}=0$.
Subcase (ii.2): $i$ is non-pendant vertex. Then, again by Remark 2.12, $(AB)_{ik}=0$ for any $k\neq i$. Thus, $b=\tilde{b_{i}}=0$, completing the proof. ∎
@@
3. Necessary and Sufficient Condition of $A$ such that $D(A),D(A^{\#})\in\mathbb{D}$
Before we proceed, let us provide two examples. First we consider a matrix
@@ 3. Necessary and Sufficient Condition of $A$ such that $D(A),D(A^{\#})\in\mathbb{D}$
Since the non-pendant vertices 2, 4, 5 are not adjacent to any pendant vertex in $D(A^{\#})$, $D(A^{\#})\notin\mathbb{D}$. On the other hand, for a matrix
@@
$D(B)$ is a star tree and belongs to $\mathbb{D}$. Since $B^{\#}=B$, $D(B^{\#})$ also belongs to $\mathbb{D}$. The two examples above serve a good motivation for our main result (Theorem 3.5) in this section.
The first result in this section shows that the $ij$-th entry of the group inverse of a real matrix $A$ is non-zero when $i,j$ are maximally matchable in $D(A)\in\mathbb{D}$.
@@
By Proposition 2.5, the length of the alternating cycle chain from $i$ to $j$ is at most three, since $D(A)\in\mathbb{D}$.
Case (i): The length of the alternating cycle chain is one, so that $(i,j,i)$ is a pendant cycle. Without loss of generality, let $i$ be the pendant vertex. Let $\{i_{1}(=i),i_{2},\ldots,i_{s}\}$ be the set of pendant vertices which have $j$, the non-pendant vertex, as a common neighbour. Then, $\mathbb{M}=\cup_{m=1}^{s}\mathbb{M}(i_{m},j),$ a disjoint union. Now,
$\Delta_{A}$ $=\sum_{m=1}^{s}\sum_{M\in\mathbb{M}(i_{m},j)}\eta(M)$
$=\sum_{m=1}^{s}a_{i_{m}j}a_{ji_{m}}\sum_{M\in\mathbb{M}(i_{m},j)}\beta_{\overline{i_{m},j}}(M)$
$=\sum_{m=1}^{s}a_{i_{m}j}a_{ji_{m}}\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)$
$=\Big{(}\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\Big{)}\sum_{m=1}^{s}a_{i_{m}j}a_{ji_{m}}.$
Since $\Delta_{A}\neq 0$, $\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\neq 0$ and so, $\alpha_{ij}=\frac{\beta_{ij}}{\Delta_{A}}\Big{(}\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\Big{)}\neq 0$.
Case (ii): The length of the alternating cycle chain is three. Let the unique alternating cycle chain from $i$ to $j$ be $((i,q,i),(q,z,q),(z,j,z))$. So, $i,j$ are pendant while $q,z$ are non-pendant vertices. Let $\{i_{1}(=i),i_{2},\ldots,i_{s}\}$ and $\{j_{1}(=j),j_{2},\ldots,j_{p}\}$ be the set of all pendant vertices having $q$ and $p$ as a common neighbour, respectively. Then, $\mathbb{M}=\cup_{m=1}^{s}\cup_{l=1}^{p}\mathbb{M}(i_{m},j_{l})$, again a mutually disjoint union. Now, we obtain
$\Delta_{A}$ $=\sum_{m=1}^{s}\sum_{l=1}^{p}\sum_{M\in\mathbb{M}(i_{m},j_{l})}\eta(M)$
$=\sum_{m=1}^{s}\sum_{l=1}^{p}(a_{i_{m}q}a_{qi_{m}})(a_{j_{l}z}a_{zj_{l}})\sum_{M\in\mathbb{M}(i_{m},j_{l})}\beta_{\overline{i_{m},j_{l}}}(M)$
$=\sum_{m=1}^{s}\sum_{l=1}^{p}(a_{i_{m}q}a_{qi_{m}})(a_{j_{l}z}a_{zj_{l}})\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)$
$=\Big{(}\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\Big{)}\sum_{m=1}^{s}\sum_{l=1}^{p}(a_{i_{m}q}a_{qi_{m}})(a_{j_{l}z}a_{zj_{l}}).$
Again, since $\Delta_{A}\neq 0$, $\sum_{M\in\mathbb{M}(i,j)}\beta_{\overline{i,j}}(M)\neq 0$ and so, $\alpha_{ij}\neq 0$. ∎
@@
In the following result, we show that when $D(A)\in\mathbb{D}$ for a real matrix $A$, $D(A^{\#})$ is a simple symmetric digraph.
@@
Let $A$ be an $n\times n$ real matrix such that $D(A)\in\mathbb{D}$. Assume that $D(A)$ is strongly connected and $\Delta_{A}\neq 0$. Then, $D(A^{\#})$ is a strongly connected simple symmetric directed graph.
@@
Since $D(A)$ is strongly connected, $A$ is an irreducible matrix. Then by [18, Lemma 2.4], $A^{\#}$ is irreducible. So, $D(A^{\#})$ is strongly connected. Let $A^{\#}=(\alpha_{ij})$. Then $\mu_{ii}=0$ implies $\alpha_{ii}=0$ and so, $D(A^{\#})$ is simple. For any $i$ and $j$, suppose $\alpha_{ij}\neq 0$. Then, by Theorem 2.8, $i,j$ are maximally matchable. Clearly, $j,i$ are also maximally matchable and so, from Corollary 3.1, $\alpha_{ji}\neq 0$. Thus, $A^{\#}$ is a combinatorially symmetric matrix and $D(A^{\#})$ is a symmetric directed graph. ∎
Recall that a simple symmetric directed graph is said to be corona graph if each non-pendant vertex is adjacent to exactly one pendant vertex.
@@
If $D(A)$ is a star tree graph, then $D(A^{\#})$ is a star tree graph.
$(ii)$ If $D(A)$ is a corona graph, then $D(A^{\#})$ is a corona graph.
@@
$(i)$ Let the vertex set of $D(A)$ be $\{1,2,\ldots,n\}$. Without loss of generality, let the center vertex of $D(A)$ be $n$. Then, the only non zero entries in $A$ are $a_{in},a_{ni}$ for all $11\leq i\leq n-$ 1. On the other hand, the maximum matchings of $D(A)$ are $M_{i}=\{(n,i,n)\}$ and so, the only maximum matchable pair of vertices are $n,i$ for all $11\leq i\leq n-$ 1. Let $A^{\#}=(\alpha_{ij})$. Then, by Theorem 2.8 and Corollary 3.1, the only non zero entries of $A^{\#}$ are $\alpha_{in},\alpha_{ni}$ for all $11\leq i\leq n-$ 1. Thus, $D(A^{\#})$ is a star tree graph.
$(ii)$ If $D(A)$ is a corona graph, then $n$ is even. Let the vertex set of $D(A)$ be $\{1,2,\ldots,\frac{n}{2},\frac{n}{2}+1,\ldots n\}$. Without loss of generality, let the set of non-pendant vertices and pendant vertices of $D(A)$ be $\{1,2,\ldots,\frac{n}{2}\}$ and $\{\frac{n}{2}+1,\frac{n}{2}+2,\ldots n\}$, respectively, where vertex $k$ is adjacent to vertex $\frac{n}{2}+k$ for all $1\leq k\leq\frac{n}{2}$. Then, the matrix $A$ is of the form 2.1, where $F$ and $G$ both are diagonal matrices of order $\frac{n}{2}$ with non-zero entries in the diagonal. Also, $A$ is invertible and

### v3

@@ 2. Graphical description for the group inverses of matrices with simple symmetric digraphs
Let $D\in\mathcal{D}$ have $k$ non-pendant vertices. Then, the number of 2-cycles in a maximum matching is always $k$. Note that, every non-pendant vertex is matched in any maximum matching of $D$, and both the endpoints of a length three alternating cycle chain are pendant vertices and a length one alternating cycle chain is nothing but a pendant cycle.
Proposition 2.3.
Let $D\in\mathcal{D}$ be a strongly connected digraph. We then have the following:
$(i)$ The length of any alternating cycle chain is at most three.
$(ii)$ Any alternating cycle chain between two vertices is unique.
Proof.
$(ii)$ Let $C$ be an alternating cycle chain from vertex $i$ to vertex $j$ in $D$. Then, by first part, the length of $C$ is at most three. If the length is one then $C$ is simply a pendant cycle. So, it is unique. Now, let us consider the case when the length of $C$ is three. Since pendant cycles are the only maximum matching cycles (corresponding to any maximum matching), the initial and the terminal vertices of $C$ should be pendant vertices. So, $i$ and $j$ are pendant vertices. Let $\tilde{C}$ be another alternating cycle chain from $i$ to $j$. Since $i$ and $j$ are pendant vertices, the initial and the terminal 2-cycles of $C$ and $\tilde{C}$ must be the same. Here, $\tilde{C}$ is an alternating cycle chain, so its length should be odd. $C$ and $\tilde{C}$ are two different alternating cycle chains with the condition that they have the same initial and terminal 2-cycles. This means that the length of the cycle chain $\tilde{C}$ should be at least five, a contradiction to the fact that the length of any alternating cycle chain in $D$ is at most three (as we proved in part (i)). Therefore, $C$ and $\tilde{C}$ coincide. ∎
Remark 2.4.
For a tree digraph, there is a unique cycle chain between two arbitrary vertices. A simple symmetric digraph $D(A)\in\mathcal{D}$ may not have a unique cycle chain between two arbitrary vertices. If a cycle chain between two vertices is an alternating cycle chain, it is always a unique alternating cycle chain between them in $D(A)$.
Let $A$ be an $n\times n$ real matrix such that $D(A)\in\mathcal{D}$. Then recall a blockwise description for $A^{\#}$ from [21]. Suppose $D(A)$ has $k$ non-pendant vertices. For $i=1,2,\ldots,k$, let $x_{i}$ and $y_{i}$ be (column) vectors of length $r_{i}\in\mathbb{N}$ such that every coordinate is nonzero $(\sum_{i=1}^{k}r_{i}=n-k)$. Set
Then $A$ can be written permutationally similar to a matrix of the form
$\begin{pmatrix}E&F\\ G&0\\ \end{pmatrix},$ (2.1)
where $E=(e_{ij})\in\mathbb{R}^{k\times k}$. Set $\alpha_{i}=x_{i}^{T}y_{i}$ for all $i\in\{1,2,\ldots,k\}$. Let $W=(W_{ij})$ be the $k\times k$ block matrix, where $W_{ij}=(\frac{e_{ij}}{\alpha_{i}\alpha_{j}})y_{i}x_{j}^{T}$, an $r_{i}\times r_{j}$ matrix. Then, by [21, Theorem 2.5], the group inverse of matrices of the form 2.1 has the $2\times 2$ block form
$\begin{pmatrix}0&~{}Y\\ Z&-W\\ \end{pmatrix},$ (2.2)
For a square matrix $A$, $A^{\#}$ may not always exist. For instance, if $A=\begin{pmatrix}0&1\\ 0&0\\ \end{pmatrix},$ then $A^{\#}$ does not exist. In particular, it is easy to show that the group inverse does not exist for any nilpotent matrix. The following proposition provides a necessary and sufficient condition for the existence of the matrix $A^{\#}$ when $D(A)\in\mathcal{D}$, that is the sum of all maximum matchings in $D(A)$ is nonzero.
Proposition 2.5.
Let $A$ be an $n\times n$ real matrix such that $D(A)\in\mathcal{D}$. Assume that $D(A)$ is strongly connected. Then $A^{\#}$ exists if and only if $\Delta_{A}\neq 0$.
Proof.
$A$ has the form 2.1 since $D(A)\in\mathcal{D}$. Now, set $\alpha_{i}=x_{i}^{T}y_{i}$ for all $i\in\{1,2,\ldots,k\}$. For each $i$, the non-pendant vertex $i$ is adjacent to $r_{i}$ pendant vertices and $\alpha_{i}$ is the sum of cycle products of $r_{i}$ pendant cycles incident $i$. Note that
By [28, Theorem 3.1], $A^{\#}$ exists if and only if $rank(F)=rank(FG)$. Here, $rank(F)=k$ which implies $A^{\#}$ exists if and only if $\alpha_{i}\neq 0$ for all $i\in\{1,2,\ldots,k\}$. By Proposition 2.1, all the maximum matching cycles are pendant cycles and a maximum matching of $D(A)$ has a set of $k$ pendant cycles incident to $k$ non-pendant vertices (the total number of maximum matching is $\prod_{i=1}^{k}r_{i}$). So, $\Delta_{A}=\prod_{i=1}^{k}\alpha_{i}$ and it follows that $A^{\#}$ exists if and only if $\Delta_{A}\neq 0$. ∎
Remark 2.6.
Let $A=(a_{ij})$ be an $n\times n$ real matrix such that $D(A)\in\mathcal{D}$. Let $\{i_{1},i_{2},\ldots,i_{s+1}\}$ be a set of all pendant vertices with a common neighbour $q$. Then,
@@
In order to prove Theorem 1.4, we need three lemmas, which we prove now.
@@
Let $A=(a_{ij})$ be an $n\times n$ real matrix such that $D(A)\in\mathcal{D}$. Let $D(A)$ be a strongly connected digraph and $\Delta_{A}\neq 0$. Let $B=(b_{ij})$ be the matrix given by $b_{ij}=\frac{\mu_{ij}}{\Delta_{A}},1\leq i,j\leq n$. Then, $AB=BA$.
@@
Now, we consider the fact $i\neq j$ and prove 2.3 by considering four cases.
@@
Subcase (iii.1): $i$ is adjacent to $j$. then, the left hand side of 2.3 is zero and the right hand side equal to
@@
Subcase (iii.2.a): $q$ is adjacent to $j$. Since non-pendant cycles do not belong to any maximum matching, again, the left hand side of 2.3 is zero, and the right hand side can be split as 2.4. Let $\mathbb{M}_{j_{l}}(i,q)\subseteq\mathbb{M}(i,q)$ be the set of maximum matching containing pendant cycle $(j,j_{l},j)$. Then $\mathbb{M}(i,q)=\cup_{l=1}^{s}\mathbb{M}_{j_{l}}(i,q)$, a mutually disjoint union. Now, the second sum of 2.4 is
@@
So, the right hand side of 2.3 is also zero.
Subcase (iii.2.b): $q$ is not adjacent to $j$. Since $q$ and $j$ both are non-pendant vertices, the left hand side of equation 2.3 is zero. After splitting the right hand side as 2.4, in the first sum, the length of any cycle chain from vertex $i$ to $j_{l}$ must be at least four for any $l\in\{1,2,\ldots,s\}$ and so, $\mu_{ij_{j}}$ is zero. Next, as $m$ is a non-pendant vertex and since it is not adjacent to $i$, the second sum is also zero, showing that the right hand side equals zero.
@@
Then, the rest of the calculations can be done by giving the same arguments as in the proof of [22, Theorem 2.8]. ∎
@@
Then, again the rest of the proof same as the proof of [22, Theorem 2.7]. ∎
Proof of Theorem 1.4.
It follows from Lemma 2.7, 2.10 and 2.11. ∎
@@
3. Necessary and sufficient condition of a real matrix $A$ such that $D(A),D(A^{\#})\in\mathcal{D}$
Before we proceed, let us provide an example.
Example 3.1.
First we consider a matrix
@@ 3. Necessary and sufficient condition of a real matrix $A$ such that $D(A),D(A^{\#})\in\mathcal{D}$
Since the non-pendant vertices 2, 4 and 5 are not adjacent to any pendant vertex in $D(A^{\#})$, $D(A^{\#})\notin\mathcal{D}$. On the other hand, for a matrix
@@
$D(B)$ is a star tree and belongs to $\mathcal{D}$. Since $B^{\#}=B$, $D(B^{\#})$ also belongs to $\mathcal{D}$.
The two examples above serve a good motivation for our main result (Theorem 1.5) in this section.
The first result in this section shows that the $ij$-th entry of the group inverse of a real matrix $A$ with $D(A)\in\mathcal{D}$ is non-zero when $i,j$ are maximally matchable in $D(A)$.
@@
The proof follows from the proof of [22, Corollary 2.12]. ∎
@@
In the following result, we show that when $D(A)\in\mathcal{D}$ for a real matrix $A$, $D(A^{\#})$ is a simple symmetric digraph.
@@
Let $A$ be an $n\times n$ real matrix such that $D(A)\in\mathcal{D}$. Assume that $D(A)$ is strongly connected and $\Delta_{A}\neq 0$. Then, $D(A^{\#})$ is a strongly connected simple symmetric digraph.
@@
Since $D(A)$ is strongly connected, $A$ is an irreducible matrix. Then by [17, Lemma 2.4], $A^{\#}$ is irreducible. So, $D(A^{\#})$ is strongly connected. Let $A^{\#}=(\alpha_{ij})$. Then $\mu_{ii}=0$ implies $\alpha_{ii}=0$ and so, $D(A^{\#})$ is simple. For any $i$ and $j$, suppose $\alpha_{ij}\neq 0$. Then, by Theorem 1.4, $i,j$ are maximally matchable. Clearly, $j,i$ are also maximally matchable and so, from Corollary 3.2, $\alpha_{ji}\neq 0$. Thus, $A^{\#}$ is a combinatorially symmetric matrix and $D(A^{\#})$ is a simple symmetric digraph. ∎
Recall that a simple symmetric digraph is said to be corona digraph if each non-pendant vertex is adjacent to exactly one pendant vertex.
@@
If $D(A)$ is a star tree digraph, then $D(A^{\#})$ is a star tree digraph.
$(ii)$ If $D(A)$ is a corona digraph, then $D(A^{\#})$ is a corona digraph.
@@
$(i)$ Let the vertex set of $D(A)$ be $\{1,2,\ldots,n\}$. Without loss of generality, let the center vertex of $D(A)$ be $n$. Then, the only non zero entries in $A$ are $a_{in},a_{ni}$ for all $11\leq i\leq n-$ 1. On the other hand, the maximum matchings of $D(A)$ are $M_{i}=\{(n,i,n)\}$ and so, the only maximum matchable pair of vertices are $n,i$ for all $11\leq i\leq n-$ 1. Let $A^{\#}=(\alpha_{ij})$. Then, by Theorem 1.4 and Corollary 3.2, the only non zero entries of $A^{\#}$ are $\alpha_{in},\alpha_{ni}$ for all $11\leq i\leq n-$ 1. Thus, $D(A^{\#})$ is a star tree digraph.
$(ii)$ If $D(A)$ is a corona digraph, then $n$ is even. Let the vertex set of $D(A)$ be $\{1,2,\ldots,\frac{n}{2},\frac{n}{2}+1,\ldots n\}$. Without loss of generality, let the set of non-pendant vertices and pendant vertices of $D(A)$ be $\{1,2,\ldots,\frac{n}{2}\}$ and $\{\frac{n}{2}+1,\frac{n}{2}+2,\ldots n\}$, respectively, where vertex $k$ is adjacent to vertex $\frac{n}{2}+k$ for all $1\leq k\leq\frac{n}{2}$. Then, the matrix $A$ is of the form 2.1, where $F$ and $G$ both are diagonal matrices of order $\frac{n}{2}$ with non-zero entries in the diagonal. Also, $A$ is invertible and